OpenStudy (anonymous):

Here is a question from my Homework (Intro to College Algebra): "Show that \(\left( \mathbb Z_{2^n} \right)^*\) is not cyclic for any \(n \ge 3\) (this is the multiplicative group consisting of the nonzero elements of \(\mathbb Z_{2^n}\)). [Hint: cyclic groups have unique subgroups of any order dividing the order of the group.]" Note: \(\mathbb Z_{2^n} = \{0, 1, 2, \ldots, 2^n-1\}\) PLEASE DON'T POST THE SOLUTION. I am quite frankly a bit confused with the problem itself. Instead, I would like clarification as to why this group would be cyclic for \(n < 3\)... it doesn't make sense

OpenStudy (anonymous):

that i can answer i think if \(n=1\) you get \(\mathbb{Z}_2\) which is clearly cyclic

OpenStudy (anonymous):

and if \(n=2\) you get \(\mathbb{Z}_4\) also cyclic

OpenStudy (anonymous):

actually now that i think about it, what exactly is \(\mathbb{Z_n}^*\) ?

OpenStudy (anonymous):

oh nvm i see it is multiplicative group of non zero elements of \(\mathbb{Z}_n\) so \(\mathbb{Z_2}^*=\{1\}\)

OpenStudy (anonymous):

now i have a question how is \(\{1,2,3\} \) the non zero elements of \(\mathbb{Z}_4\) a group under multiplication? what is the inverse of 2?

OpenStudy (anonymous):

I have no idea haha...that's why the problem is confusing me... it seems blatantly obvious that \((\mathbb Z_{2^n})^*\) even for any \(n\) wouldn't be cyclic anyway...